Circular No. 6828

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

V1333 AQUILAE J. H. Swank and E. Smith, Goddard Space Flight Center (GSFC);

and A. M. Levine and R. Remillard, Massachusetts Institute of Technology (MIT), on behalf on the RXTE ASM team at GSFC and MIT, report that the low-mass x-ray binary Aql X-1 (= V1333 Aql) has started another x-ray outburst. In real-time results from the RXTE All Sky Monitor, the daily average flux levels in the band 2–12 keV are: Feb. 27, ∼ 22 mCrab; 28, 36; Mar. 1, 59; 2, 100; 3, 127. Pointed observations with the PCA began on Mar. 3 and confirm the ASM results. Observations at other wavelengths are encouraged.

XTE J1858+034 F. E. Marshall, Goddard Space Flight Center; D. Chakrabarty, Mas-

sachusetts Institute of Technology; and M. H. Finger, Marshall Space Flight Center, report an improved position for XTE J1858+034 (IAUC 6826) of α = 18^{h}58^{m}.6, δ = +3^{o}21′ (equinox 2000.0), with an estimated 90-percent confidence error radius of 2′.5, based on repeated scans across the source with the PCA on RXTE. The flux (2–10 keV) on Feb. 28.9 UT was 4 mCrab.

COMETS 55P/TEMPEL-TUTTLE, 103P/HARTLEY 2, 69P/TAYLOR

D. K. Lynch and R. W. Russell, Aerospace Corporation; and M. Sitko, University of Cincinnati, report observations of comets 55P, 103P and 69P using BASS on the NASA Infrared Telescope Facility. The heliocentric distances of the comets were 1.03, 1.24, and 2.01 AU, respectively, with 55P approaching and 103P and 69P receding from the sun. Comet 55P/Tempel-Tuttle had an N magnitude of 4.0 ± 0.12 on Feb. 8.25 UT. The spectrum showed little evidence of silicate emission and was well fit with a single 330-K blackbody between 3 and 13 μm, some 60 K above the equilibrium blackbody temperature of 274 K; observations on Feb. 9.25 showed this comet to be unchanged from the previous night within the formal uncertainties. Comet 103P/Hartley 2 had an N magnitude of 4.8 ± 0.2 on Feb. 9.25 and showed a weak silicate emission feature rising ∼ 15 percent above the continuum; the 8–13-μm color temperature was 320 K, ∼ 70 K above the blackbody equilibrium temperature of 250 K for a dust particle at the computed heliocentric distance. Comet 69P/Taylor had an N magnitude of 4.5 ± 0.20 on Feb. 9.4 and showed a small infrared excess indicative of silicate emission; the 8–13-μm color temperature was 250 K, again ∼ 50 K above the blackbody equilibrium temperature.

1998 March 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green